# ON EQUICONTINUITY OF MAPPINGS WITH VARIABLE DEFINITION DOMAINS

### v1

@@ 1 Introduction
This paper is devoted to the study of mappings satisfying upper distortion estimates of the modulus of families of paths. We consider the case when definition domains of mappings may be different. The main subject of this study is distortion estimates for the distance under mappings. For fixed domains, this problem has been repeatedly studied by many authors, including both homeomorphisms and mappings with a branching. In particular, we could cite papers on this matter, see, e.g., [Cr1][Cr2], [MRSY], [RSY1][RSY2], [RS1], [Sev1] and [SevSkv]. We note that we draw heavily on the ideas of Suvorov [Suv], who considered sequences of domains and the theory of prime ends for them. At the same time, the present paper is not, generally speaking, concerned with the theory of prime ends.
@@
Given $p\geqslant 1,$ a number $\delta>0,$ a fixed domain $D_{0}\subset{\mathbb{R}}^{n},$ $n\geqslant 2,$ a sequence of domains $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ the kernel of which is $D_{0},$ a non-degenerate continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{F}_{Q,A,p,\delta}(D_{0},\mathfrak{D})$ some a family of ring $Q$-homeomorphisms $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ such that $h(f_{m}(A))\geqslant\delta$ and $h(\overline{{\mathbb{R}}^{n}}\setminus f_{m}(D_{m}))\geqslant\delta.$
@@
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$
In addition, assume that $p=n$ and:
5) there is $r>0$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m});$
6) for each point $x_{0}\in D_{0}$ and for every $0<r_{1}<r_{2}<r_{0}:=\sup\limits_{x\in D_{0}}|x-x_{0}|$ there is a set $E_{1}\subset[r_{1},r_{2}]$ of a positive linear Lebesgue measure such that the function $Q$ is integrable with respect to $\mathcal{H}^{n-1}$ over the spheres $S(x_{0},r)$ for every $r\in E_{1}.$ (In particular, the latter holds if $Q$ is a locally integrable function).
Then $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ and some $\delta_{1}>0.$
@@
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$ Besides that, if $p=n$ and conditions 5)–6) of Theorem 1 hold, then for any continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ such that $h(f_{m}(A))\geqslant\delta$ for any $m=1,2,\ldots,$ there exists $\delta_{1}>0$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$
@@ 2 Preliminaries
On the other hand, observe that $f_{m}(\Gamma_{m})=\Gamma(C_{m},f_{m}(A_{m}),D_{m}^{\,\prime}).$ By the hypothesis of the lemma, $h(f_{m}(A_{m}))\geqslant\delta$ for all $m\in{\mathbb{N}}.$ Therefore, due to (2.6), $h(f_{m}(A_{m}))\geqslant\delta_{1}$ and $h(C_{m})\geqslant\delta_{1},$ where $\delta_{1}:=\min\{\delta,a/2\}.$ Using the fact that the domains $D_{m}^{\,\prime}:=f_{m}(D)$ are equi-uniform with respect to the $p$-modulus, we conclude that there exists $\sigma>0$ such that
@@
3 On image of the continuum under homeomorphisms
@@
Given numbers $\delta,r,N>0,$ a family $\mathfrak{D}$ of domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{I}_{Q,\delta,N}(\mathfrak{D})$ a family of all open discrete and closed mappings $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ defined in some $D=D_{f}\in\mathfrak{D}$ and satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{\mathfrak{D}}$ for all $0<r_{1}<r_{2}<\infty$ for which there exists a non-degenerate continuum $A=A_{f}$ and balls $\overline{B(x^{1}_{f},r^{1}_{f})},\ldots,\overline{B(x^{N}_{f},r^{N}_{f})},$ $x^{i}_{f}\in A_{f},$ $1\leqslant i\leqslant N,$ such that
$(1)\qquad A_{f}\subset\bigcup\limits_{i=1}^{N}\overline{B(x^{i}_{f},r^{i}_{f})}\subset D_{f}\,$,
$(2)\qquad h(\overline{B(x^{i}_{f},r^{i}_{f})},\partial D_{f})\geqslant\delta\,$,
$(3)\qquad h(f(A_{f}))\geqslant\delta$
and satisfying the condition: $h(E)\geqslant r$ whenever $E$ is any component of $\partial f(D_{f}).$ The following statement was proved for a fixed domain, see [RS2, Lemma 3]. Now we extend it for mappings defined on the family of domains.
Lemma 3.2.  Let $f_{m}\in\mathfrak{I}_{Q,\delta,N}(\mathfrak{D}).$ Assume that, for each point $x_{0}\in D_{0}$ and for every $0<r_{1}<r_{2}<r_{0}:=\sup\limits_{x\in D}|x-x_{0}|$ there is a set $E_{1}\subset[r_{1},r_{2}]$ of a positive linear Lebesgue measure such that the function $Q$ is integrable with respect to $\mathcal{H}^{n-1}$ over the spheres $S(x_{0},r)$ for every $r\in E_{1}.$ Then there exists $\delta_{1}>0$ such that $h(f_{m}(A_{m}),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots,$ where $A_{m}:=A_{f_{m}},$ $D_{m}:=D_{f_{m}}.$
The statement of Lemma 3 remains true if, instead of the condition regarding the integrability of the function $Q$ over spheres with respect to some set $E_{1}$ is replaced by a simpler condition: $Q\in L_{\rm loc}^{1}({\mathbb{R}}^{n}).$
@@
Proof.   Since $D_{m}$ is a domain in ${\mathbb{R}}^{n},$ $m=1,2,\ldots,$ $\partial D_{m}\neq\varnothing.$ In addition, since $C(\partial D_{m},f_{m})\subset\partial f_{m}(D_{m})$ whenever $f_{m}$ is open, discrete and closed (see [Vu, Theorem 3.3]), $\partial f_{m}(D_{m})\neq\varnothing,$ as well. Thus, the quantity $h(f_{m}(A_{m}),\partial f_{m}(D_{m}))$ is well-defined, so the formulation of the lemma is correct.
@@
Let us prove Lemma 3 by the contradiction, partially using the approach of the proof of Lemma 4.1 in [SevSkv]. Suppose that the conclusion of the lemma is not true. Then for each $k\in{\mathbb{N}}$ there is some number $m=m_{k}$ such that $h(f_{m_{k}}(A_{m_{k}}),\partial f_{m_{k}}(D_{m_{k}}))<1/k,$ $A_{m}:=A_{f_{m}}.$ In order not to complicate the notation, we will further assume that $h(f_{m}(A_{m}),\partial f_{m}(D_{m}))<1/m,$ $m=1,2,\ldots.$ Note that the set $f_{m}(A_{m})$ is compact as a continuous image of a compact set $A_{m}\subset D_{m}$ under the mapping $f_{m}.$ In this case, there are elements $x_{m}\in f_{m}(A_{m})$ and $y_{m}\in\partial f_{m}(D_{m})$ such that $h(f_{m}(A_{m}),\partial f_{m}(D_{m}))=h(x_{m},y_{m})<1/m.$ Let $E_{m}$ be a component of $\partial f_{m}(D_{m})$ containing $y_{m}.$ Due to the compactness of $\overline{{\mathbb{R}}^{n}},$ we may assume that $y_{m}\rightarrow y_{0}$ as $m\rightarrow\infty;$ then also $x_{m}\rightarrow y_{0}$ as $m\rightarrow\infty.$
@@
Put $P>0$ and $U=B_{h}(y_{0},r_{0})=\{y\in\overline{{\mathbb{R}}^{n}}:h(y,y_{0})<r_{0}\},$ where $2r_{0}:=\min\{r/2,\delta/2\},$ where $r$ and $\delta$ are numbers from the condition of the lemma. Observe that $E_{m}\cap U\neq\varnothing\neq E_{m}\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $y_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $y_{m}\in E_{m};$ besides that $h(E_{m})\geqslant r>r/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Observe that, $E_{m}$ is closed (see [Ku, Theorem 1.III.5.46]). Now, $E_{m}$ is a continuum, so that $E_{m}\cap\partial U\neq\varnothing$ by Proposition 3. Similarly, $f_{m}(A_{m})\cap U\neq\varnothing\neq f_{m}(A_{m})\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $x_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $x_{m}\in f_{m}(A_{m});$ besides that $h(f_{m}(A_{m}))\geqslant\delta>\delta/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Since $f_{m}(A_{m})$ is a continuum, $f_{m}(A_{m})\cap\partial U\neq\varnothing$ by Proposition 3. By the proving above,
$E_{m}\cap\partial U\neq\varnothing\neq f_{m}(A_{m})\cap\partial U$ (3.1)
@@
$E_{m}\cap\partial V\neq\varnothing\neq f_{m}(A_{m})\cap\partial V$
@@
$M(\Gamma(f_{m}(A_{m}),E_{m},\overline{{\mathbb{R}}^{n}}))>P$ (3.3)

### v2

# arXiv:2602.14013v2[math.CV] 01 Mar 2026

@@ 1 Introduction
This paper is devoted to the study of mappings satisfying upper distortion estimates of the modulus of families of paths. We consider the case when definition domains of mappings may be different. The main subject of this study is distortion estimates for the distance under mappings. For fixed domains, this problem has been repeatedly studied by many authors, including both homeomorphisms and mappings with a branching. In particular, we could cite papers on this matter, see, e.g., [Cr1][Cr2], [MRSY], [M1][M3], [RSY1][RSY2], [RS], [Sev1] and [SevSkv]. We note that we draw heavily on the ideas of Suvorov [Suv], who considered sequences of domains and the theory of prime ends for them. At the same time, the present paper is not, generally speaking, concerned with the theory of prime ends.
@@
Given $p\geqslant 1,$ a number $\delta>0,$ a fixed domain $D_{0}\subset{\mathbb{R}}^{n},$ $n\geqslant 2,$ a sequence of domains $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ the kernel of which is $D_{0},$ a non-degenerate continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{F}_{Q,A,p,\delta}(D_{0},\mathfrak{D})$ some a family of ring $Q$-homeomorphisms $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ for all $0<r_{1}<r_{2}<\infty$ such that $h(f_{m}(A))\geqslant\delta$ and $h(\overline{{\mathbb{R}}^{n}}\setminus f_{m}(D_{m}))\geqslant\delta.$
@@
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$ In addition, assume that
5) $p=n$ and there is $r>0$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m}).$
Then $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ and some $\delta_{1}>0.$ Moreover, if $A_{0}$ is some another compactum in $D_{0},$ then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
@@
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$ Besides that, if the condition 5) of Theorem 1 holds, then for any continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ such that $h(f_{m}(A))\geqslant\delta$ for any $m=1,2,\ldots,$ there exists $\delta_{1}>0$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$
Results similar to Theorems 1 and 1 also hold for domains with prime ends. We now formulate these results.
The next definitions due to Caratheodory [Car]; cf. [KR]. Let $\omega$ be an open set in ${\mathbb{R}}^{k}$, $k=1,\ldots,n-1$. A continuous mapping $\sigma\colon\omega\rightarrow{\mathbb{R}}^{n}$ is called a $k$-dimensional surface in ${\mathbb{R}}^{n}$. A surface is an arbitrary (n1)(n-1)-dimensional surface $\sigma$ in ${\mathbb{R}}^{n}.$ A surface $\sigma$ is called a Jordan surface, if $\sigma(x)\neq\sigma(y)$ for $x\neq y$. In the following, we will use $\sigma$ instead of $\sigma(\omega)\subset{\mathbb{R}}^{n},$ $\overline{\sigma}$ instead of $\overline{\sigma(\omega)}$ and $\partial\sigma$ instead of $\overline{\sigma(\omega)}\setminus\sigma(\omega).$ A Jordan surface $\sigma\colon\omega\rightarrow D$ is called a cut of $D$, if $\sigma$ separates $D,$ that is $D\setminus\sigma$ has more than one component, $\partial\sigma\cap D=\varnothing$ and $\partial\sigma\cap\partial D\neq\varnothing$.
A sequence of cuts $\sigma_{1},\sigma_{2},\ldots,\sigma_{m},\ldots$ in $D$ is called a chain, if:
(i) the set $\sigma_{m+1}$ is contained in exactly one component $d_{m}$ of the set $D\setminus\sigma_{m},$ wherein $\sigma_{m-1}\subset D\setminus(\sigma_{m}\cup d_{m})$; (ii) $\bigcap\limits_{m=1}^{\infty}\,d_{m}=\varnothing.$
Two chains of cuts $\{\sigma_{m}\}$ and $\{\sigma_{k}^{\,\prime}\}$ are called equivalent, if for each $m=1,2,\ldots$ the domain $d_{m}$ contains all the domains $d_{k}^{\,\prime}$, except for a finite number, and for each $k=1,2,\ldots$ the domain $d_{k}^{\,\prime}$ also contains all domains $d_{m},$ except for a finite number.
The end of the domain $D$ is the class of equivalent chains of cuts in $D$. Let $K$ be the end of $D$ in ${\mathbb{R}}^{n}$, then the set $I(K)=\bigcap\limits_{m=1}\limits^{\infty}\overline{d_{m}}$ is called the impression of $K$. Following [Na], we say that the end $K$ is a prime end, if $K$ contains a chain of cuts $\{\sigma_{m}\}$ such that
$\lim\limits_{m\rightarrow\infty}M(\Gamma(C,\sigma_{m},D))=0$
for some continuum $C$ in $D.$ The following notation is used: the set of prime ends corresponding to the domain $D,$ is denoted by $E_{D},$ and the completion of the domain $D$ by its prime ends is denoted $\overline{D}_{P}.$
Consider the following definition, which goes back to Näkki [Na], cf. [KR]. The boundary of a domain $D$ in ${\mathbb{R}}^{n}$ is said to be locally quasiconformal if every $x_{0}\in\partial D$ has a neighborhood $U$ that admits a quasiconformal mapping $\varphi$ onto the unit ball ${\mathbb{B}}^{n}\subset{\mathbb{R}}^{n}$ such that $\varphi(\partial D\cap U)$ is the intersection of ${\mathbb{B}}^{n}$ and a coordinate hyperplane. The sequence of cuts $\sigma_{m},$ $m=1,2,\ldots,$ is called quasiconformally regular, if $\overline{\sigma_{m}}\cap\overline{\sigma_{m+1}}=\varnothing$ for $m\in{\mathbb{N}}$ and, in addition, $d(\sigma_{m})\rightarrow 0$ as $m\rightarrow\infty.$ If the end $K$ contains at least one quasiconformally regular chain, then $K$ will be called quasiconformally regular. We say that a bounded domain $D$ in ${\mathbb{R}}^{n}$ is quasiconformally regular, if $D$ can be quasiconformally mapped to a domain with a locally quasiconformal boundary whose closure is a compact in ${\mathbb{R}}^{n},$ and, besides that, every prime end in $D$ is quasiconformally regular. Note that space $\overline{D}_{P}=D\cup E_{D}$ is metric, which can be demonstrated as follows. If $g:D_{0}\rightarrow D$ is a quasiconformal mapping of a domain $D_{0}$ with a locally quasiconformal boundary onto some domain $D,$ then for $x,y\in\overline{D}_{P}$ we put:
$\rho(x,y):=|g^{\,-1}(x)-g^{\,-1}(y)|\,$, (1.7)
where the element $g^{\,-1}(x),$ $x\in E_{D},$ is to be understood as some (single) boundary point of $D_{0}.$ The specified boundary point is unique and well-defined by [IS, Theorem 2.1, Remark 2.1], cf. [Na, Theorem 4.1]. It is easy to verify that $\rho$ in (1.7) is a metric on $\overline{D}_{P}.$ If $g_{*}$ is another quasiconformal mapping of a domain $D_{*}$ with locally quasiconformal boundary onto $D$, then the corresponding metric $\rho_{*}(p_{1},p_{2})=|{\widetilde{g_{*}}}^{-1}(p_{1})-{\widetilde{g_{*}}}^{-1}(p_{2})|$ generates the same convergence and, consequently, the same topology in $\overline{D}_{P}$ as $\rho_{0}$ because $g_{0}\circ g_{*}^{-1}$ is a quasiconformal mapping of $D_{*}$ and $D_{0}$, which extends, by Theorem 4.1 in [Na], to a homeomorphism between $\overline{D_{*}}$ and $\overline{D_{0}}$. In the sequel, this topology in $\overline{D}_{P}$ will be called the topology of prime ends; the continuity of mappings $F\colon\overline{D}_{P}\rightarrow\overline{D^{\,\prime}}_{P}$ will be understood relative to this topology.
We say that the boundary of $D$ is weakly flat at a point $x_{0}\in\partial D$ if, for every number $P>0$ and every neighborhood $U$ of the point $x_{0},$ there is a neighborhood $V\subset U$ such that $M_{\alpha}(\Gamma(E,F,D))\geqslant P$ for all continua $E$ and $F$ in $D$ intersecting $\partial U$ and $\partial V.$ We say that the boundary $\partial D$ is weakly flat if the corresponding property holds at every point of the boundary.
Let $D_{m},$ $m=1,2,\ldots,$ be a sequence of domains which converges to a kernel $D_{0}.$ Then $D_{m}$ will be called regular with respect to $D_{0}$ in the terms of prime ends, if $D_{m}\subset D_{0}$ for all $m\in{\mathbb{N}}$ and, for every $P_{0}\in E_{D_{0}}$ there is a sequence $\sigma_{k},$ $k=1,2,\ldots,$ with the following condition: if $d_{k}$ is a domain in $P_{0}$ then there is $M=M(k)$ such that $d_{k}\cap D_{m}$ is a non-empty connected set for every $m\geqslant M(k).$ The following result holds.
1) the sequence of domains $D_{m}$ is regular with respect to $D_{0}$ in terms of prime ends;
2) for every $m\in{\mathbb{N}},$ a domain $D_{m}$ is quasiconformally regular; $D_{0}$ is also quasiconformally regular;
4) at least one of two following conditions hold: $Q$ has a finite mean oscillation in $\overline{D_{0}},$ or (1.6) holds for every $x_{0}\in\overline{D_{0}}$ and some $\beta(x_{0})>0.$
Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D}$ in terms of prime ends, i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $\rho(x,y)<\delta,$ $m\in{\mathbb{N}}$ and $\rho$ is one of the metric in $\overline{D_{0}}_{P}.$
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}$ with respect $\rho.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}_{P}\rightarrow\overline{{\mathbb{R}}^{n}}.$ In addition, assume that
5) $p=n$ and there is $r>0$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m}).$
Then $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ and some $\delta_{1}>0.$ Moreover, if $A_{0}$ is some another compactum in $D_{0},$ then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D}$ with respect to prime ends, i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $\rho(x,y)<\delta,$ $m\in{\mathbb{N}}$ and $\rho$ is one of the metric in $\overline{D_{0}}_{P}.$
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$ Besides that, if the condition 5) of Theorem 1 holds, then for any continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ such that $h(f_{m}(A))\geqslant\delta$ for any $m=1,2,\ldots,$ there exists $\delta_{1}>0$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$
@@ 2 Preliminaries
Due to [MRSY, Section 3] we say that a boundary $D$ is called strongly accessible with respect to $p$-modulus at $x_{0}\in\partial D,$ if for any neighborhood $U$ of the point $x_{0}\in\partial D$ there is a neighborhood $V\subset U$ of this point, a compactum $F\subset D$ and a number $\delta>0$ such that $M_{p}(\Gamma(E,F,D))\geqslant\delta$ for any continua $E\subset D$ such that $E\cap\partial U\neq\varnothing\neq E\cap\partial V.$ The boundary of a domain $D$ is called strongly accessible with respect to $p$-modulus, if this is true for any $x_{0}\in\partial D.$ When $p=n,$ prefix ’’relative to $p$-modulus’’ is omitted.
Remark 2.1. Uniform domains have strongly accessible boundaries, see e.g. [SevSkv, Remark 1]. The inverse statement is also true at least for $p=n,$ see [Sev3, Lemma 4.2].
@@
On the other hand, observe that $f_{m}(\Gamma_{m})=\Gamma(C_{m},f_{m}(A_{m}),D_{m}^{\,\prime}).$ By the hypothesis of the lemma, $h(f_{m}(A_{m}))\geqslant\delta$ for all $m\in{\mathbb{N}}.$ Therefore, due to (2.6), $h(f_{m}(A_{m}))\geqslant\delta_{1}$ and $h(C_{m})\geqslant\delta_{1},$ where $\delta_{1}:=\min\{\delta,a/2\}.$ Using the fact that the domains $D_{m}^{\,\prime}:=f_{m}(D_{m})$ are equi-uniform with respect to the $p$-modulus, we conclude that there exists $\sigma>0$ such that
@@
3 On image of the continuum under open discrete and closed mappings
@@
Lemma 3.2.  Let $D_{0}\subset{\mathbb{R}}^{n},$ be a domain in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and let $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ be a sequence of domains the kernel of which is $D_{0}.$ Let $A$ be a non-degenerate continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ and let $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$ be a Lebesgue measurable function. In addition, let $f_{m}:D_{m}\rightarrow{\mathbb{R}}$ be a sequence of open discrete and closed mappings satisfying relations (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ for all $0<r_{1}<r_{2}<d(x_{0},\partial D_{m})$ such that $h(f_{m}(A))\geqslant\delta>0$ for any $m=1,2,\ldots$ and $p=n$ for some $\delta>0.$ Assume that, $Q$ has a finite mean oscillation in $A,$ or the relation (1.6) holds for every $x_{0}\in A$ and some $\beta(x_{0})>0.$ If there is $r>0$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m}),$ then there exists $\delta_{1}>0$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$
Proof.   Observe that, $h(A,\partial D_{m})\geqslant r_{*}>0$ for some $r_{*}>0,$ and all $m\in{\mathbb{N}}$ except finite numbers. Otherwise, there are sequences $x_{m}\in A$ and $y_{m}\in\partial D_{m}$ with $h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty.$ We may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ Obviously, $x_{0}\in A$ by the closeness of $A.$ Now, there is $\varepsilon_{0}>0$ such that $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}$ and $\overline{B(x_{0},\varepsilon_{0})}$ belongs to all $D_{m}$ except of finite numbers. However, the latter contradicts the condition $y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty$ while $y_{m}\in\partial D_{m}.$ Indeed, in this case $\overline{B(x_{0},\varepsilon_{0})}\cap\partial D_{m}\neq\varnothing$ for infinitely many $m\in{\mathbb{N}}.$
We firstly consider the case when the relation (1.6) holds. Since for any $0<\varepsilon_{1}<\widetilde{\varepsilon_{1}}<\min\{\beta(x_{0}),\frac{1}{2}r_{*}\}$ we have that
$\int\limits_{\varepsilon_{1}}^{\widetilde{\varepsilon_{1}}}\frac{dt}{tq_{x_{0}}^{\,\prime\,\frac{1}{n-1}}(t)}\leqslant\int\limits_{\varepsilon_{1}}^{\widetilde{\varepsilon_{1}}}\frac{dt}{t}<\infty\,$,
the relation (1.6) implies that, given $0<\widetilde{\varepsilon_{1}}<r_{*}\leqslant h(A,\partial D_{m})$ we may find $\varepsilon_{1}=\varepsilon_{1}(x_{0})$ such that
$\int\limits_{\varepsilon_{1}}^{\widetilde{\varepsilon_{1}}}\frac{dt}{tq_{x_{0}}^{\,\prime\,\frac{1}{n-1}}(t)}>1\,.$ (3.1)
We cover $A$ with the balls $B(x_{0},\varepsilon_{1}),$ $x_{0}\in A,$ where $\varepsilon_{1}$ as above. By Heine-Borel-Lebesgue lemma we may find a finite subcover $B(x_{1},\varepsilon_{1}),$ $B(x_{2},\varepsilon_{2}),\ldots,B(x_{N},\varepsilon_{N})$ such that
$A\subset\bigcup\limits_{i=1}^{n}B(x_{i},\varepsilon_{i})\,,\qquad A\subset\bigcup\limits_{i=1}^{n}B(x_{i},2\varepsilon_{i})\,.$ (3.2)
@@
By the choice of $\varepsilon_{i}>0,$ $\varepsilon_{i}<\min\{\beta(x_{0}),\frac{1}{2}r_{*}\},$ we have that $B(x_{i},2\varepsilon_{i})\subset D_{m}.$ Since $D_{m}$ is a domain in ${\mathbb{R}}^{n},$ $m=1,2,\ldots,$ $\partial D_{m}\neq\varnothing.$ In addition, since $C(\partial D_{m},f_{m})\subset\partial f_{m}(D_{m})$ whenever $f_{m}$ is open, discrete and closed (see [Vu, Theorem 3.3]), $\partial f_{m}(D_{m})\neq\varnothing,$ as well. Thus, the quantity $h(f_{m}(A),\partial f_{m}(D_{m}))$ is well-defined, so the formulation of the lemma is correct.
@@
Let us prove Lemma 3 by the contradiction, partially using the approach of the proof of Lemma 4.1 in [SevSkv]. Suppose that the conclusion of the lemma is not true. Then for each $k\in{\mathbb{N}}$ there is some number $m=m_{k}$ such that $h(f_{m_{k}}(A),\partial f_{m_{k}}(D_{m_{k}}))<1/k.$ In order not to complicate the notation, we will further assume that $h(f_{m}(A),\partial f_{m}(D_{m}))<1/m,$ $m=1,2,\ldots.$ Note that the set $f_{m}(A)$ is compact as a continuous image of a compact set $A\subset D_{m}$ under the mapping $f_{m}.$ In this case, there are elements $x_{m}\in f_{m}(A)$ and $y_{m}\in\partial f_{m}(D_{m})$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))=h(x_{m},y_{m})<1/m.$ Let $E_{m}$ be a component of $\partial f_{m}(D_{m})$ containing $y_{m}.$ Due to the compactness of $\overline{{\mathbb{R}}^{n}},$ we may assume that $y_{m}\rightarrow y_{0}$ as $m\rightarrow\infty;$ then also $x_{m}\rightarrow y_{0}$ as $m\rightarrow\infty.$
@@
Put $P>0$ and $U=B_{h}(y_{0},r_{0})=\{y\in\overline{{\mathbb{R}}^{n}}:h(y,y_{0})<r_{0}\},$ where $2r_{0}:=\min\{r/2,\delta/2\},$ where $r$ and $\delta$ are numbers from the condition of the lemma. Observe that $E_{m}\cap U\neq\varnothing\neq E_{m}\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $y_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $y_{m}\in E_{m};$ besides that $h(E_{m})\geqslant r>r/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Observe that, $E_{m}$ is closed (see [Ku, Theorem 1.III.5.46]). Now, $E_{m}$ is a continuum, so that $E_{m}\cap\partial U\neq\varnothing$ by Proposition 3. Similarly, $f_{m}(A)\cap U\neq\varnothing\neq f_{m}(A)\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $x_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $x_{m}\in f_{m}(A);$ besides that $h(f_{m}(A))\geqslant\delta>\delta/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Since $f_{m}(A)$ is a continuum, $f_{m}(A)\cap\partial U\neq\varnothing$ by Proposition 3. By the proving above,
$E_{m}\cap\partial U\neq\varnothing\neq f_{m}(A)\cap\partial U$ (3.3)
@@
$E_{m}\cap\partial V\neq\varnothing\neq f_{m}(A)\cap\partial V$
@@
$M(\Gamma(f_{m}(A),E_{m},\overline{{\mathbb{R}}^{n}}))>P$ (3.5)